f={(1,2),(2,3),(3,3),(4,5)},g={(1,1)(2,3),(3,2),(4,8)}
Find (f*g)(2)=

Answer by MathLover1(20855)


: f={(1,2),(2,3),(3,3),(4,5)},
g={(1,1)(2,3),(3,2),(4,8)}
(F *G)(x) = F(x) * G(x)
(F *G)(1) = F(1) *G(1) = 2 *1 =2
(F *G)(2) = F(2) *G(2) = 3*3 = 9
(F *G)(3) = F(3) *G(3) = 3* 2 = 6
(F *G)(4) = F(4) *G(4) = 5*8=40
(F *G)(x) ={(1,2),(2,9),(3,6),(4,40)}
(F *G)(2) =9